<note important>Achtung! R1 TRANSFER IN PROGRESS! MONITORING REQUIRED!</note> R1 monitoring instructions: operations:monitoring.ivs.etransfer

Hobart 12m:

Disk VSN: HOB+0126/32000

Data volume at beginning: 26535.894 GB

  • 09:30 UT - Disc pos ~60 GB behind, lost 30 in the last 2 1/2 hours. (Ellen)
  • 10:35 UT - Swapped to module GSFC+022. Data on HOB+0126: 31966.008 GB. First scan 134-1035. (Ellen)

Katherine 12m:

Disk VSN: HOB+0128/32000

Data volume at beginning: 13239.838 GB

Yarragadee 12m:

Disk VSN: HOB+0115/16000

Data volume at beginning: 6620.624 GB

Parkes 64m:

Experiment started OK

06:00 UT - Handed over Parkes to Stacy (vk)